FULL ANALYSIS OF THREE-STATE-VIBRONIC COUPLING IN THE PYRAZOLYL-d_{3} RADICAL

\underline{TAKATOSHI ICHINO}, ADAM J. GIANOLA, and W. CARL LINEBERGER, JILA and Department of Chemistry and Biochemistry, University of Colorado, Boulder, CO 80309; JOHN F. STANTON, Institute for Theoretical Chemistry and Department of Chemistry and Biochemistry, The University of Texas at Austin, Austin, TX 78712.

The 351.1 nm photoelectron spectrum of pyrazolide-d_{3} anion has been measured. The photoelectron angular distributions indicate the presence of nearly-degenerate electronic states of the pyrazolyl-d_{3} radical. Equation-of-motion coupled-cluster calculations (EOMIP-CCSD) have been performed to study low-lying states of the radical. The calculations lead us to conclude that three electronic states, energetically close to each other, are accessed in the photodetachment process. Strong interactions of the pseudo-Jahn-Teller type within each pair of the three states are evident in the calculations for the radical at the anion geometry. Model diabatic potentials of the three states have been constructed around the anion geometry in terms of the anion reduced normal coordinates up to second-order. Parameters of the model potentials have been derived from the EOMIP-CCSD. Simulation of the spectrum based on the model Hamiltonian has been performed, treating all the vibronic interactions among the three states simultaneously. The simulation reproduces the fine structure of the observed spectrum very well, revealing complicated non-adiabatic effects on the spectrum. The ground state of pyrazolyl-d_{3} is ^{2}A_{2}, and the electron affinity is 2.934 ± 0.006 eV. The first excited state is ^{2}B_{1}, with a term energy of 32 ± 1 meV. While the stationary points of the X̃^{2}A_{2} and Ã^{2}B_{1} states are minima, that of B̃^{2}B_{2} state is a saddle point as a result of the pseudo-Jahn-Teller interactions with the other two states. The topology of the model potential energy surfaces is discussed. Supported by NSF and AFOSR.